### v1

# DRIFT: Decompose, Retrieve, Illustrate, then Formalize Theorems

Abstract
Automating the formalization of mathematical statements for theorem proving remains a major challenge for Large Language Models (LLMs). LLMs struggle to identify and utilize the prerequisite mathematical knowledge and its corresponding formal representation in languages like Lean. Current retrieval-augmented autoformalization methods query external libraries using the informal statement directly, but overlook a fundamental limitation: informal mathematical statements are often complex and offer limited context on the underlying math concepts. To address this, we introduce Drift, a novel framework that enables LLMs to decompose informal mathematical statements into smaller, more tractable “sub-components”. This facilitates targeted retrieval of premises from mathematical libraries such as Mathlib. Additionally, Drift retrieves illustrative theorems to help models use premises more effectively in formalization tasks. We evaluate Drift across diverse benchmarks (ProofNet, ConNF, and MiniF2F-test) and find that it consistently improves premise retrieval, nearly doubling the F1 score compared to the DPR baseline on ProofNet. Notably, Drift demonstrates strong performance on the out-of-distribution ConNF benchmark, with BEq+@10 improvements of 37.14% and 42.25% using GPT-4.1 and DeepSeek-V3.1, respectively. Our analysis shows that retrieval effectiveness in mathematical autoformalization depends heavily on model-specific knowledge boundaries, highlighting the need for adaptive retrieval strategies aligned with each model’s capabilities.
1 Introduction
Autoformalization is formulated as a translation task that aims to translate natural language mathematical descriptions into machine-verifiable statements written in formal languages, such as Coq (Barras et al., 1997), Isabelle (Paulson, 1994), and Lean (De Moura et al., 2015). Previous work has shown that accurate autoformalization is a critical step towards developing automated theorem proving systems (Lin et al., 2025b; Chen et al., 2025; Xin et al., 2024; Lin et al., 2025c), and ultimately assist mathematicians in new discoveries (Gouëzel & Shchur, 2019; Leang et al., 2025). Despite recent progress of Large Language Models (LLMs) in informal mathematical reasoning (Ahn et al., 2024; Setlur et al., 2024; Luong & Lockhart, 2025), formal reasoning presents distinct challenges. The strict syntax and necessity for precise alignment between informal concepts and formal definitions mean that even powerful pretrained models often fail at autoformalization tasks off-the-shelf (Wu et al., 2025).
Although synthetic data generation could enable the fine-tuning of LLMs for autoformalization (Jiang et al., 2023; Lin et al., 2025a; Wang et al., 2024; Ying et al., 2024), the knowledge cut-off issue raised by updating formal libraries like Mathlib (Mathlib Community, 2020) makes finetuned models prone to hallucinating non-existent formal objects that have been renamed, reorganized or deprecated (Baanen et al., 2025). Early retrieval-augmented methods addressed this by retrieving similar theorems from external libraries to provide useful syntactic structure and compositional examples. However, their practical utility as exemplars has been limited by the retrieval methods used to find them. These methods often lack task-specific training data and rely on general-purpose techniques like keyword searching (Agrawal et al., 2022), k-NN (Azerbayev et al., 2023), or pretrained dense retrievers (Zhang et al., 2024). Liu et al. (2025) introduced dependency retrieval task and developed the RAutoformalizer (RAuto) framework. Similar to the premise selection for proof generation (Yang et al., 2023), this paradigm enables the training of specialized retrievers to identify the exact definitions that formal statements require. However, this new approach created a key trade-off: focusing on individual components meant losing the valuable context provided by full theorem statements. Based on this, we identify two main limitations in the current approach to retrieval-augmented autoformalization: Informal statements are often dense and multifaceted. This underlying complexity of queries makes them suboptimal as direct queries for retrieving the precise, atomic definitions required for formalization. Additionally, even when the correct formal definitions are retrieved, models often lack the knowledge of contextual usage required to correctly structure and integrate them into the formal statement.
In information retrieval, query enhancement techniques like query expansion (Chan et al., 2024), pseudo-document generation (Gao et al., 2023), and neural query rewriting (Wang et al., 2025) have demonstrated effectiveness of reformulating queries to provide more semantic information. Query decomposition has proven particularly useful for multi-hop question answering, as it matches the granularity of the dense query statements with the indexed documents (Ammann et al., 2025). In addition to retrieving correct premises, providing rich context like exemplar proofs can guide proof generation effectively (He et al., 2024; Thakur et al., 2024; Thompson et al., 2025). Despite advances in adjacent domains, these techniques have not yet been systematically applied to dependency retrieval for autoformalization, which still relies on monolithic queries and provides context-free definitions.
We propose Drift, a novel framework depicted in Figure 1, that enhances retrieval-augmented autoformalization by adapting query decomposition and context augmentation to address unique challenges of theorem autoformalization. To address complex queries, we adapt query decomposition to break down complex informal statements into a series of simpler, atomic sub-queries. Each sub-query targets a single mathematical concept, transforming a multifaceted retrieval problem into focused, precise searches. Second, to illustrate the usage of the dependencies, we contextualize the retrieved definitions with demonstrative theorems, giving the model concrete examples of syntax and application patterns.
Contributions. 1) We introduce Drift 111The code and models are available at https://github.com/huawei-noah/DRIFT. (Decompose, Retrieve, Illustrate, then Formalize Theorems), a decomposition-driven retrieval augmented formalization framework that breaks down informal mathematical statements into atomic sub-queries and contextualizes retrieved premises with demonstrative theorems, transforming monolithic retrieval into a process aligned with the dependency structure of formal mathematics. 2) Our experiments establish new state-of-the-art in dependency retrieval and autoformalization on ProofNet and ConNF across models, with exceptional performance on out-of-distribution ConNF benchmark. 3) Through systematic analysis, we establish that the utility of retrieved dependencies is conditioned on the gap between a model’s parametric knowledge and the statement complexity. These insights reveal critical design considerations for retrieval-augmented systems and point toward the necessity of adaptive strategies that can assess when external knowledge genuinely complements model capabilities.
2 Related Work
Retrieval-augmented Autoformalization. Early retrieval-augmented autoformalization methods retrieved similar theorems as few-shot examples: ProofNet (Azerbayev et al., 2023) employs k-NN search, while MS-RAG (Zhang et al., 2024) uses informal-to-informal retrieval with iterative refinement. LTRAG (Hu et al., 2025) retrieves thought-guided theorem pairs for neuro-symbolic formalization. Liu et al. (2025) established this paradigm by introducing “dependency retrieval”, a premise selection task specialized for autoformalization: given an informal statement, retrieve the precise set of formal objects and definitions that are required for its autoformalization from a library (e.g. Mathlib). While RAutoformalizer (RAuto) (Liu et al., 2025) demonstrated improvements over non-retrieval methods, evaluation revealed a significant gap compared to oracle systems with ground-truth dependencies. CRANF (Lu et al., 2025) attempts conceptual mapping between abstraction levels. Critically, all existing approaches treat complex statements as monolithic queries, failing to identify distinct mathematical concepts within them.
Query Decomposition and Enhancement. Query enhancement has proven effective across retrieval tasks. Query2Doc (Wang et al., 2023) and HyDE (Gao et al., 2023) generate pseudo-documents to expand semantic coverage. SemanticSearch (Gao et al., 2024) augment the informal statement by prompting an LLM to translate the query into a detailed statement with both informal and formal expressions. However, they only evaluated with similar theorem retrieval but not on the downstream formalization. More relevant to our work, query decomposition has shown success in multi-hop question answering (Ammann et al., 2025), where breaking complex questions into sub-queries improves retrieval of distinct information aspects. Zhao et al. (2023) and Jiang et al. (2022) applied similar decompositions to theorem proving, showing the effectiveness of divide and conquer in formal math. Despite these successes, no prior work has applied decomposition to informal mathematical statements for autoformalization.
3 Methodology
We introduce Drift (Decompose, Retrieve, Illustrate, then Formalize Theorems), a novel four-stage method designed to address the two main limitations of previous retrieval-augmented formalization methods, 1) handling the complexity of informal statements and 2) the lack of demonstrative examples for retrieved formal objects. As a first step, an LLM decomposes the informal statement into a set of smaller, granular sub-queries (§3.1). These queries then guide the retrieval of dependent premises from a formal library such as Mathlib (§3.2). In a subsequent, bottom-up illustration step, we find theorems that utilize the retrieved premises, providing in-context demonstrations of their application (§3.3). Finally, the LLM formalizes the original statement, conditioned on all retrieved premises and theorems (§3.4). This process is visualized in Figure 1.
3.1 Decompose
Standard retrieval methods often treat complex informal statements as monolithic queries and simply embed the entire statement. This approach disregards the rich semantic structure within a statement, which may contain multiple distinct mathematical concepts. Their complexity means that a statement could be under-specified, ambiguous and/or simply too information-dense to be used as is. Compressing the entire statement’s meaning into a single dense vector creates a representational bottleneck, risking the loss of nuanced details and focusing the retrieval on only the most salient concepts. We hypothesize that by decomposing an informal statement into its constituent concepts, we can perform a more granular and accurate retrieval for each concept.
To this end, Drift begins with a Decompose module (panel ① in Figure 1), which is implemented as an LLM tasked with breaking down an informal statement () into a set of structured sub-queries, , see Equation 1. While the decomposer could be a finetuned model, we leverage in-context learning (ICL) with off-the-shelf LLMs in this study.
Each sub-query is designed to isolate a single mathematical concept. As illustrated in Figure 1, the component is a natural language phrase describing the concept (e.g., “A prime number of the form is a prime that leaves remainder 3 when divided by 4”) while is a predicted formal representation for that concept (e.g., “ : with the conditions Nat.Prime and % , where % denotes the modulo operation on natural numbers.”). Appending this predicted formal name serves a dual purpose. First, it probes the LLM’s parametric knowledge, providing a syntactic “anchor” for the concept. Second, it allows the retriever to jointly leverage the semantics of the natural language phrase and the syntactic cues from to identify the correct premise even in the presence of minor inaccuracies in the predicted formal representations.
3.2 Retrieve
The Retrieve module is designed to identify dependent premises from the library that correspond to the concepts isolated in each sub-query. This one-to-one mapping is visualized in the ② panel of Figure 1, which shows each sub-query () being linked to a formal object (e.g., PNat.Prime). To accomplish this, we implement a dense passage retriever (Karpukhin et al., 2020) using a BGE-M3 (Chen et al., 2024) encoder (), finetuned on the dependency prediction task as introduced by Liu et al. (2025). This training objective aligns the formal dependencies with their informal statement, thereby making semantic similarity a strong proxy for logical dependency. We achieve this by encoding queries and formal library objects into a shared -dimensional embedding space. The vector representations for all formal objects are pre-computed in an offline step and stored in an efficient search index. At inference, each sub-query is encoded into a vector and the closest dependent premise is identified by finding the library object that maximizes the cosine similarity score, with the query vector. This is formally defined as:
where . The final set of dependent premises, , is formed by aggregating the top-1 result from each sub-query and removing duplicates: .
3.3 Illustrate
Retrieving useful formal definitions is a necessary first step, however, it is not sufficient for successful autoformalization. For instance, a retrieved definition like def ZMod : → Type tells the model what the concept is, but provides no further guidance on its practical application such as the syntax for declaring a variable of that type (a : ZMod p). This gap between definition and usage is a primary source of syntactic and structural errors in LLM-generated formal statements.
The Illustrate module is designed to bridge this gap by providing examples of formal object usage, visualized in Figure 1 (panel ③). Given a premise like ZMod from the “Retrieve” step, the module selects illustrative theorems, e.g. “Theorem 2” which demonstrates the correct application of ZMod. The module takes the set of retrieved premises , and a budget , as input, and outputs a small set of illustrative theorems , where . The selection process is a greedy algorithm designed to maximize the coverage of the input premises as follows. First, we compile a candidate set of all theorems in library that utilize at least one of the retrieved premises .
In order to ensure relevance and provide a deterministic tie-breaker, we pre-sort this candidate set by the semantic similarity of each theorem’s informal statement to the original informal statement . The cosine similarity is computed using vectors from the same encoder as the retrieval stage: .
The final set of illustrative theorems is built iteratively. The process begins by initializing an empty set of selected theorems and an empty set of covered premises . At each step , we select the theorem that provides the maximal marginal gain by including the most new premises in not previously covered:
where is the set of premises used in theorem . After selecting , the sets are updated for the next iteration: and . The process terminates when either the budget of theorems is reached or when no remaining candidate theorem can offer additional coverage (i.e., the marginal gain is zero). The final set from the last iteration .
3.4 Formalize Theorems
The final Drift step is Formalize Theorems, shown in panel ④ of Figure 1. The formalizer assembles the previously gathered definitions and theorem demonstrations into a comprehensive prompt and generates the final formal statement. The formalization module is designed to be flexible and can be implemented with either a finetuned model or a general-purpose LLM guided with ICL. The formalizer compiles information in the following logical order: , where denotes the concatenation operator, is a formalization instruction (details in Section A.3.2), are the retrieved premises, are the illustrative theorems and is the original informal statement. Conditioned on this comprehensive prompt, the formalizer then generates the final output of the Drift framework, the formal statement .
4 Experimental Setup
4.1 Benchmarks
We evaluate Drift on three distinct autoformalization benchmarks to test its in-distribution, self-contained and out-of-distribution performance. While the experiments are conducted in Lean, our framework is language-agnostic and adaptable to other formal systems with structured libraries.
ProofNet (In-Distribution). We use the ProofNet benchmark (Azerbayev et al., 2023) for in-distribution evaluation. Its 374 theorems, sourced from undergraduate mathematics textbooks, are integrated with the Mathlib library and require an average of 3.39 dependent premises from over 243k formal objects (including 139k theorems). This benchmark tests the framework’s primary function, which is to effectively retrieve and utilize dependent premises with demonstration from a large-scale, in-distribution knowledge base.
MiniF2F (Self-Contained). We use the MiniF2F-test set (Zheng et al., 2021) to evaluate the framework on self-contained problems. This benchmark consists of 224 Olympiad-style theorems with a notably low average of just 0.43 dependencies from Mathlib.222We removed duplicate and low-quality pairs that do not compile with our Lean version (4.18.0). MiniF2F-test serves as a boundary condition, testing the model’s core formalization capabilities and its robustness against potentially distracting context when retrieval is not strictly necessary.
ConNF (Out-of-Distribution). The OOD challenge refers to evaluation scenarios where neither the retrieval model nor the formalization model has been exposed to the formal objects used in the test data.333The knowledge cutoff dates of the models we used are June 2024, March 2025 and July 2025, for GPT-4.1, Claude-Opus-4, and DeepSeek-3.1, respectively. Though there is a risk of exposure to the formalizer which cannot be controlled or confirmed, from our zero-shot results on ConNF, it is hypothesised that the models have not been extensively trained on this benchmark. We therefore evaluate on ConNF, a benchmark curated by Liu et al. (2025) through topological informalization to test OOD generalization. This benchmark is based on the con-nf library and is not integrated with Mathlib.444The con-nf library is available at https://github.com/leanprover-community/con-nf. It formalizes a consistency proof for Quine’s New Foundations (Quine, 1951), established by Holmes & Wilshaw (2015) and contains 961 research-level theorems requiring an average of 3.92 premises from its distinct library of 1,348 formal objects. ConNF provides a rigorous test of Drift’s ability to generalize to a novel mathematical domain (Zhang et al., 2024).
4.2 Baselines
We evaluate our proposed framework against three key baselines representing zero-shot, state-of-the-art retrieval, and oracle-level performance. We note that concurrent work, CRAMF (Lu et al., 2025), was not publicly available for comparison at the time of our experiments. The no retrieval (zero-shot) baseline establishes a performance floor. The autoformalization model receives only the informal statement as input, without access to any retrieved context. DPR (RAuto) represents the current state-of-the-art. We compare our method to the dependency retrieval module from the RAutoformalizer framework (Liu et al., 2025), which is a finetuned dense passage retriever (DPR) (Karpukhin et al., 2020).555Model available at https://huggingface.co/purewhite42/dependency_retriever_f. The top-5 retrieved premises are provided to the formalizer model as augmented context. The oracle* (retrieval) setting provides an approximate upper bound for retrieval. The model is provided with ground-truth dependencies () for each problem, simulating a ’perfect’ retriever, as defined by Liu et al. (2025). We mark this setting with an asterisk (*) to denote that this oracle is in fact imperfect because the provided dependencies are not necessarily optimal or exhaustive for formalization and do not necessarily lead to the best autoformalization performance as we discuss in Section 5, i.e. some of our settings are actually outperforming this imperfect oracle.
4.3 Implementation Details
In this study, we evaluate Drift as a lightweight prompting strategy, leveraging the in-context learning capabilities of instruction-following models such as DeepSeek-3.1 (DeepSeek-AI, 2024), GPT-4.1 (OpenAI, 2025), and Claude-Opus-4 (Anthropic, 2025) to demonstrate Drift’s broad applicability without introducing task-specific biases from finetuning. While we use the same model for both decomposition and formalization for consistency, these modules are independent and could be replaced with specialized finetuned models.
Decompose. We construct a few-shot prompt using five expertly-verified examples from the Putnam benchmark (Tsoukalas et al., 2024) to decompose the informal statements; full details in Section A.1.1 and Section A.3.1. Each decomposed sub-query consists of a natural language description of a concept and its formal representation retrieved from parametric knowledge, as detailed in Section 3.1. Through this, we instruct the LLMs to decompose problems.
Retrieve. The retriever model is a dense passage retriever666Implementation details in Section A.1.2 (DPR) finetuned on Mathlib data to map informal queries to their formal dependencies (Liu et al., 2025). We pre-compute embeddings for all formal declarations in the relevant libraries (Mathlib for ProofNet and MiniF2F-test; con-nf for ConNF). This training setup establishes ConNF as an OOD benchmark for the retrieval module. For the baseline, we retrieve the top-5 premises based on the entire informal statement.
Illustrate. In order to demonstrate premise usage in real contexts, we select up to exemplar theorems using the greedy coverage algorithm described in Section 3.3.
Formalize Theorems. As described in Section 3.4, the formalization prompt combines the original informal statement with the retrieved premises and illustrative theorems. Each premise is presented with its full name, formal declaration, and source code. Each illustrative theorem is included as a pair of its informal and formal statements. This demonstrates both the informal-to-formal alignment and the concrete application of the premises within a theorem instance. To evaluate pass@k, we generate 10 formalizations for each problem. The prompt context, including the retrieved premises and the illustrative theorems, remains consistent across all 10 runs.
4.4 Evaluation Metrics
We ablate the evaluation of Drift in two stages: a) the intrinsic performance of dependency retrieval plus the selection of illustrative theorems, and b) the extrinsic performance of autoformalization.
Intrinsic Dependency Retrieval. The effectiveness of the decomposition is measured by its impact on the dependency retrieval task as the quality of the decomposed sub-queries directly impacts the relevance of the retrieved premises. We measure the quality of the retrieved premises against the oracle* dependencies using Precision, Recall, and their harmonic mean, F1-score.
Formalization Correctness. For extrinsic evaluation, we use Typecheck (TC) and BEq+. Typecheck measures syntactic correctness, indicating the percentage of the generated statements that are valid and can pass the compiler’s type checker (Lu et al., 2024; Azerbayev et al., 2023; Liu et al., 2025). For semantic correctness, we use BEq+ (Poiroux et al., 2025), a symbolic metric that measures the logical equivalence between a predicted formal statement and the ground-truth reference by using deterministic proof tactics to bidirectionally prove that each statement can be transformed into the other. For each metric, we assess performance using pass@1 and pass@10, where pass@ indicates that at least one of independent generations was successful.
5 Results and Discussion
5.1 Dependency Retrieval
| Benchmark | Decomposer | Precision | Recall | F1 |
| ProofNet | - | 11.55 | 17.03 | 13.77 |
| Claude-Opus-4 | 23.02 | 34.70 | 27.68 | |
| GPT-4.1 | 21.71 | 34.46 | 26.64 | |
| DeepSeek-3.1 | 24.38 | 30.28 | 27.01 | |
| MiniF2F-test | - | 0.36 | 4.12 | 0.66 |
| Claude-Opus-4 | 2.08 | 23.71 | 3.83 | |
| GPT-4.1 | 1.42 | 15.46 | 2.60 | |
| DeepSeek-3.1 | 0.98 | 9.28 | 1.78 | |
| ConNF | - | 25.12 | 32.06 | 28.17 |
| Claude-Opus-4 | 30.97 | 41.01 | 35.29 | |
| GPT-4.1 | 31.62 | 40.64 | 35.56 | |
| DeepSeek-3.1 | 34.67 | 39.39 | 36.88 |
We evaluate the effectiveness of the Decompose and Retrieve modules by looking at their impact on intrinsic performance in dependency retrieval. As detailed in Table 1, we compare Drift against a no-decomposition baseline that uses the same dense retriever but with the original informal statements as queries. This provides a direct comparison to Drift, which retrieves a similar number of premises by taking the union of the top-1 results for each of the 5.21 to 6.42 sub-queries generated by the Decompose module. The results show that decomposition provides a substantial performance improvement in both precision and recall. Averaged across all decomposer models, Drift achieves an absolute improvement of 13.34, 2.08, and 7.74 points over the baseline F1 score on the ProofNet, MiniF2F-test, and ConNF benchmarks, respectively. Regarding the choice of decomposer, we observe that while Claude-Opus-4 achieves the highest F1 scores on ProofNet (27.68%) and MiniF2F-test (3.83%), the performance variation among the LLMs is marginal. Our findings indicate that frontier LLMs are largely interchangeable for this task, as the top-performing models have a maximum F1 score difference of only 2.05% on any benchmark.
The Illustrate module proves highly effective at selecting a concise set of theorems to demonstrate premise usage. Within a maximum of only three selected theorems (), the algorithm achieves a high average premise coverage rate of 74.594.80% across all the decomposers and benchmarks.
5.2 Formalization
| Dataset | Formalizer | Retrieval | TC@1 | BEq+@1 | TC@10 | BEq+@10 |
| ProofNet | GPT-4.1 | Oracle* | 58.82 | 20.32 | 79.68 | 27.54 |
| Zero-shot | 34.22 | 9.36 | 51.60 | 13.37 | ||
| DPR (RAuto) | 51.60 | 14.71 | 73.53 | 19.25 | ||
| Drift | 55.88 | 17.38 | 77.01 | 21.93 | ||
| DeepSeek-3.1 | Oracle* | 71.12 | 21.93 | 82.09 | 27.54 | |
| Zero-shot | 60.43 | 15.51 | 71.93 | 20.32 | ||
| DPR (RAuto) | 63.37 | 17.38 | 73.53 | 19.52 | ||
| Drift | 72.73 | 18.18 | 79.41 | 20.59 | ||
| MiniF2F-test | GPT-4.1 | Oracle* | 75.45 | 23.66 | 89.29 | 30.36 |
| Zero-shot | 69.64 | 23.21 | 84.82 | 28.12 | ||
| DPR (RAuto) | 77.23 | 24.55 | 92.41 | 32.14 | ||
| Drift | 74.55 | 24.55 | 92.41 | 29.02 | ||
| DeepSeek-3.1 | Oracle* | 77.68 | 23.21 | 87.50 | 28.12 | |
| Zero-shot | 76.34 | 22.77 | 87.50 | 27.23 | ||
| DPR (RAuto) | 75.89 | 22.77(0.00) | 87.95 | 27.68 | ||
| Drift | 74.11 | 22.77(0.00) | 88.84 | 24.55 | ||
| ConNF | GPT-4.1 | Oracle* | 60.46 | 48.28 | 75.23 | 58.90 |
| Zero-shot | 7.28 | 4.47 | 11.45 | 6.76 | ||
| DPR (RAuto) | 24.56 | 15.19 | 31.95 | 20.08 | ||
| Drift | 65.76 | 54.84 | 77.00 | 62.33 | ||
| DeepSeek-3.1 | Oracle* | 57.34 | 44.22 | 71.28 | 55.15 | |
| Zero-shot | 13.42 | 8.12 | 17.59 | 11.03 | ||
| DPR (RAuto) | 21.96 | 12.90 | 28.20 | 17.07 | ||
| Drift | 60.67 | 46.72 | 71.18 | 54.21 |
On autoformalization, Drift consistently outperforms both the zero-shot baseline and the strong retrieval-augmented baseline DPR (RAuto) on ProofNet and ConNF benchmarks, both on Typecheck and BEq+ with pass@1 and pass@10 (see Table 2). Specifically, on ProofNet, GPT-4.1 with Drift achieves a BEq+ pass@10 of 21.93%, a 2.74% improvement over DPR (RAuto). This trend holds across models in our main evaluation, demonstrating that our decomposition-driven approach provides more effective context for the formalization task. A key insight from our results is that while the frontier models we tested are all highly proficient and largely interchangeable as query decomposers, this parity does not extend to the final formalization step. For instance, DeepSeek-3.1 generally outperforms GPT-4.1 in the zero-shot setting across all benchmarks, suggesting a stronger parametric knowledge for direct formalization. However, this trend reverses when retrieval is introduced, most significantly on the OOD ConNF benchmark. On ConNF, all models achieve low zero-shot BEq+ scores (10%), confirming a severe knowledge gap. Retrieval substantially improves performance, where GPT-4.1 consistently outperforms DeepSeek-3.1 across all metrics with different retrieval strategies. Drift provides a particularly substantial improvement, increasing GPT-4.1’s BEq@10 score by 55.57%, even surpasses the Oracle* baseline by 3.43%. We hypothesize this is because the Oracle* provides only necessary premises based on the reference statement, while the illustrative theorems selected with Drift provide crucial demonstrations of premise usage.
On the in-distribution ProofNet benchmark, the results are more nuanced. GPT-4.1 surpasses DeepSeek-3.1 on BEq+@10 when using Drift. This pattern suggests that GPT-4.1 is more adept at in-context synthesis, integrating and reasoning over retrieved information to construct formal statements, especially in unfamiliar domains. The distinction suggests that the two stages of our pipeline rely on distinct LLM capabilities. Decomposition leverages natural language reasoning whereas formalization demands advanced formal reasoning.
As we discussed in Section 4, the MiniF2F-test benchmark presents a distinct profile with an average of only 0.43 library dependencies. This limits the potential for retrieval-based improvements, as evidenced by the small gap between the zero-shot and Oracle* performance, e.g. a pass@10 gap of only 2.24% for GPT-4.1 and 0.89% for DeepSeek-3.1. Instead, this low-dependency regime reveals the models’ high sensitivity to the provided context, which can act as a distractor rather than an aid.
5.3 Ablation study
| Retrieval | ProofNet | MiniF2F-test | ConNF | |||
| TC@1 | BEq+@1 | TC@1 | BEq+@1 | TC@1 | BEq+@1 | |
| Drift (GPT-4.1) | 55.88 | 17.38 | 74.55 | 24.55 | 65.76 | 54.84 |
| w/o Illustrate | 56.15 (+0.27) | 14.17 (-3.21) | 76.34 (+1.79) | 24.55 ( 0.00) | 45.47 (-20.29) | 35.90 (-18.94) |
| w/o Decompose | 50.80 (-5.08) | 13.64 (-3.74) | 76.34 (+1.79) | 26.34 (+1.79) | 59.63 (-6.13) | 46.72 (-8.12) |
| w/o Retrieval | 34.22 (-21.66) | 9.36 (-8.02) | 69.64 (-4.91) | 23.21 (-1.34) | 7.28 (-58.48) | 4.47 (-50.37) |
| Drift (DeepSeek-3.1) | 72.73 | 18.18 | 74.11 | 22.77 | 60.67 | 46.72 |
| w/o Illustrate | 70.32 (-2.41) | 15.24 (-2.94) | 77.68 (+3.57) | 21.43 (-1.34) | 41.31 (-19.36) | 29.34 (-17.38) |
| w/o Decompose | 64.97 (-7.76) | 15.78 (-2.40) | 77.68 (+3.57) | 20.98 (-1.79) | 50.36 (-10.31) | 38.81 (-7.91) |
| w/o Retrieval | 60.43 (-12.30) | 15.51 (-2.67) | 76.34 (+2.23) | 22.77 ( 0.00) | 13.42 (-47.25) | 8.12 (-38.60) |
In order to isolate and measure the contribution of each component of Drift, we conducted a systematic ablation study (Table 3). As expected, removing the illustrative theorems (w/o Illustrate) significantly decreased the BEq+ score on ProofNet and ConNF, which confirms that demonstrative examples of premise usage are crucial for the formalization correctness beyond just the definitions of the formal objects. Intriguingly, additionally removing the Decompose module (w/o Decompose) does not further degrade performance and even leads to a slight recovery on ConNF and ProofNet in BEq+ score. We hypothesize this is because the baseline DPR retrieves a thematically homogeneous (though less precise) set of premises via single query retrieval, which may be less distracting than the precise but more diverse set retrieved via decomposition. This reveals a crucial synergy: the illustrative theorems act as a scaffold that helps the model navigate the diverse information from the decomposer.
The complex interaction is further highlighted by the observations on MiniF2F-test, where removing theorems improved Typecheck rate but reduced BEq+ scores that measures the logical correctness of the generated formal statement. This further supports the hypothesis that for the simpler, low-dependency problems in MiniF2F-test, adding more context can act as a distractor. Removing external context improves syntactic validity but degrades logical correctness of the generated formal statements. This sensitivity strongly motivates the need for more dynamic and adaptive retrieval strategies. Future work on agentic frameworks could selectively retrieve information, judge its utility and iterate based on the compiler feedback.
6 Conclusion
In this work, we introduced Drift, a framework that improves autoformalization by tackling two distinct challenges: the underlying complexity of queries and the lack of contextual usage. Our decomposition-driven retrieval addresses the former by breaking down the informal statement into sub-queries and conducting point-to-point retrieval of its formal dependencies. Concurrently, our theorem selector resolves the latter by providing illustrative examples to guide the utilization of retrieved premises in theorem instances. This dual approach substantially improves formalization correctness on complex and out-of-distribution benchmarks, demonstrating its effectiveness as a broadly generalizable and model-agnostic strategy. On simpler, low-dependency MiniF2F-test dataset, our method performs on-par with related methods. Our findings suggest future work to focus on dynamic and adaptive retrieval strategies, as well as on agentic frameworks that iteratively refine attempts based on compiler feedback.
Appendix A Appendix
A.1 Implementation Details
A.1.1 Few-Shot Example Generation for Decomposition
To construct a robust set of few-shot demonstrations for our decomposition module, we strategically selected five problems from the Putnam benchmark (Tsoukalas et al., 2024). These problems were chosen to ensure diversity in both their mathematical domain and the number of underlying premises required for their proofs.
We decomposed the informal statement of each selected problem into its atomic logical components using a zero-shot prompting strategy with Claude-Opus-4 in think mode, guided by a carefully engineered instruction set. To ensure the correctness and logical atomicity of these decompositions, each one was manually verified.
This curated set of examples, detailed below, provides the model with varied demonstrations for the decomposition task across number theory, algebra, analysis, and geometry.
Number Theory: putnam_1966_b2
Algebra: putnam_2000_b1
Analysis: putnam_2000_a4, putnam_2015_b1
Geometry: putnam_2003_b5
A.1.2 Retriever Finetuning Details
We finetuned the BGE-M3 retriever model (Chen et al., 2024) on the Mathlib 4.7 dataset to specialize it for dependency retrieval in formal mathematics. The fine-tuning process was executed using the FlagEmbedding library (Chen et al., 2024) on a server equipped with four 32GB GPUs. The complete set of hyperparameters used for this process, including optimizer settings and loss configuration, is provided in Table 4.
| Category | Hyperparameter | Value |
| Model & Data | model_name_or_path | bge-m3 |
| train_data | mathlib 4.7 | |
| query_max_len | 1024 | |
| passage_max_len | 1024 | |
| train_group_size | 4 | |
| sentence_pooling_method | cls | |
| Training | num_train_epochs | 1 |
| per_device_train_batch_size | 32 | |
| per_device_eval_batch_size | 4 | |
| learning_rate | ||
| warmup_ratio | 0.1 | |
| weight_decay | 0.01 | |
| repetition_penalty | 1.0 | |
| dataloader_drop_last | True | |
| even_batches | True | |
| non_blocking | False | |
| split_batches | False | |
| use_seedable_sampler | True | |
| Loss & Objective | temperature | 0.02 |
| normalize_embeddings | True | |
| negatives_cross_device | True | |
| same_benchmark_within_batch | True | |
| unified_finetuning | True | |
| kd_loss_type | m3_kd_loss | |
| Optimizer | optim | adamw_torch |
| adafactor | False | |
| adam_beta1 | 0.9 | |
| adam_beta2 | 0.999 | |
| adam_epsilon |
A.1.3 LLM Generation Parameters
For all generative tasks, sub-query generation (decomposition) and formal statement generation (formalization), we set the temperature to 0.7 for all models (GPT-4.1, DeepSeek-3.1, and Claude-Opus-4) to encourage diverse yet coherent outputs. To ensure reproducibility, single-attempt evaluations (Pass@1) used a fixed seed of 42. For multi-attempt evaluations (Pass@10), we generated ten distinct outputs by using a sequential range of seeds from 42 to 51.
A.2 Additional Results and Discussion
A.2.1 Dependency Retrieval Performance Metrics
We evaluate retrieval performance using standard precision, recall, and their harmonic mean, the F1 score. For a given retrieved set and the ground-truth set of oracle premises , precision and recall are defined as: The F1 score provides a single, balanced measure of performance by combining precision and recall: The composition of the retrieved set varies by method.
For baseline retrievers, consists of the top- premises with the highest cosine similarity to the embedding of the full informal statement. For Drift, is the union of the single best-retrieved premise for each of the decomposed sub-queries.
A.2.2 Dependency Retrieval Results of RAuto
This section presents the performance of DPR (RAuto) across both in-distribution (ProofNet, MiniF2F-test) and out-of-distribution (ConNF) benchmarks. The results, detailed in Table 5, highlight a crucial trade-off between specialization and generalization that motivates our proposed approach. When comparing the baselines, our baseline retriever substantially outperforms DPR (RAuto) on the ConNF benchmark but underperforms on ProofNet and MiniF2F-test. We hypothesize that this discrepancy arises because DPR (RAuto) may be overfitted to Mathlib-specific content.
| Dataset | Precision | Recall | F1 |
| ProofNet | 22.89 | 33.75 | 27.28 |
| MiniF2F-test | 0.63 | 7.22 | 1.15 |
| ConNF | 14.01 | 17.88 | 15.71 |
A.2.3 Autoformalization Results of Claude
Table 6 summarizes the pass@1 results of Claude-Opus-4, reporting both the Typecheck rate (TC@1) and the equivalence rate (BEq+@1) across the ProofNet, MiniF2F-test, and ConNF benchmarks. These results provide context for the performance Claude-Opus-4 discussed in the main paper.
| Dataset | Retrieval | TC@1 | BEq+@1 |
| ProofNet | Oracle* | 81.28 | 26.20 |
| Zero-shot | 68.45 | 17.65 | |
| RAuto | 75.67 | 19.25 | |
| Drift | 78.61 | 19.79 | |
| MiniF2F-test | Oracle* | 93.30 | 35.27 |
| Zero-shot | 95.09 | 31.25 | |
| RAuto | 95.98 | 30.36 | |
| Drift | 93.75 | 32.59 | |
| ConNF | Oracle* | 62.75 | 50.36 |
| Zero-shot | 13.32 | 8.53 | |
| RAuto | 26.64 | 18.52 | |
| Drift | 72.32 | 60.35 |
A.2.4 The Role of Illustrative Theorems as a Scaffold
As presented in Table 3, removing the query decomposer (w/o Decompose: reverting to the baseline DPR) does not degrade performance further. In fact, on ConNF and ProofNet, it led to a slightly recovery in the BEq+ score compared to the “w/o Illustrate” setting. We hypothesize this is due to the nature of retrieval noise. The baseline DPR, using a single query, retrieves a thematically clustered set of premises. While its precision is lower, its noise is homogeneous and may be less distracting to the LLM. Our decomposition method retrieves more diverse set of premises. While this captures more correct dependencies (higher recall and precision), the accompanying noise is also more varied. This reveals a crucial synergy: the theorem selector acts as a contextual scaffold, helping the model navigate the diverse information retrieved by the decomposer. Without this guidance, the varied noise can outweigh the benefit of improved retrieval.
A.2.5 Scaling Performance with Sampling
Across all experiments in Table 2, we observe a consistent and significant gap between pass@1 and pass@10 results. For instance, performance on ProofNet improved by an average of 27.20% across all settings and formalizer models. This large uplift underscores the potential for enhancing performance through sampling-based methods at test time. This suggests that performance could be further scaled by integrating our method into an agentic framework equipped with a verifier (e.g Typecheck correctness).
Notably, DeepSeek-3.1’s performance of pass@10 saturates more quickly than Claude-Opus-4’s on the ProofNet benchmark. Its zero-shot Pass@10 score is only 0.27% lower than its retrieval-augmented Drift score. This suggests that with sufficient sampling, the model can sometimes recover the necessary knowledge parametrically. We anticipate a similar, albeit slower, trend for the larger Claude-Opus-4 and GPT-4.1 models if the number of attempts were increased further.
A.2.6 Parametric Retrieval
| Retrieval | ProofNet | MiniF2F-test | ConNF | |||
| TC@1 | BEq+@1 | TC@1 | BEq+@1 | TC@1 | BEq+@1 | |
| Drift (GPT-4.1) | 55.88 | 17.38 | 74.55 | 24.55 | 65.76 | 54.84 |
| Parametric Retrieval | 43.85 (-12.03) | 13.64 (-3.74) | 63.84 (-10.71) | 20.09 (-4.46) | 10.41 (-55.35) | 3.64 (-51.20) |
| w/o Retrieval | 34.22 (-21.66) | 9.36 (-8.02) | 69.64 (-4.91) | 23.21 (-1.34) | 7.28 (-58.48) | 4.47 (-50.37) |
| Drift (DeepSeek-3.1) | 72.73 | 18.18 | 74.11 | 22.77 | 60.67 | 46.72 |
| Parametric Retrieval | 69.25 (-3.48) | 19.79 (+1.61) | 76.79 (+2.68) | 24.55 (+1.78) | 10.51 (-50.16) | 5.93 (-40.79) |
| w/o Retrieval | 60.43 (-12.30) | 15.51 (-2.67) | 76.34 (+2.23) | 22.77 ( 0.00) | 13.42 (-47.25) | 8.12 (-38.60) |
To distinguish between the benefits of an LLM’s internal (parametric) knowledge and external (retrieved) knowledge, we conducted a “parametric retrieval” experiment. In this setting, we prompted the formalizer models only with the decomposed sub-queries, omitting the retrieved premises and illustrative theorems. This setup probes whether the structured sub-queries alone are sufficient to guide the models to access their own latent knowledge for the formalization task.
The results in Table 7 show that external knowledge from retrieval is largely indispensable and cannot be fully replaced by the LLM’s internal knowledge alone. However, we observe a notable difference between the models. DeepSeek-3.1 demonstrates a stronger grasp of the required formal knowledge; for this model, the sub-queries appear to function as a Chain-of-Thought-style prompt, structuring its reasoning process and thereby improving formalization accuracy. This aligns with our earlier finding that DeepSeek-3.1’s zero-shot performance with sufficient sampling (pass@10) nearly saturates to its retrieval-augmented performance, suggesting it can often surface the necessary formal knowledge if properly prompted.
A.2.7 Statistics of Decomposed Sub-queries
To better understand the Decompose module and its behavior, we analyzed the number of sub-queries generated when decomposing informal statements from our three benchmarks: ProofNet, MiniF2F-test and ConNF with different LLMs as decomposer.
| Model | ProofNet | MiniF2F-test | ConNF | Model Avg. |
| Claude-Opus-4 | 5.84 | 5.59 | 6.52 | 5.98 |
| GPT-4.1 | 6.39 | 5.40 | 7.03 | 6.27 |
| DeepSeek-3.1 | 4.83 | 4.65 | 5.71 | 5.06 |
| Dataset Avg. | 5.69 | 5.21 | 6.42 | 5.77 |
The results, summarized in Table 8, show that different models produce a varying number of sub-queries. GPT-4.1 tends to generate the most detailed decompositions, with an average of 6.27 sub-queries, while DeepSeek-3.1 produces the most concise ones, averaging 5.06. Furthermore, the complexity of the dataset appears to influence the decomposition length. Statements from the ConNF benchmark, which covers frontier mathematical research, consistently required more sub-queries (6.42 on average) across all models, likely reflecting their greater conceptual density compared to the undergraduate-level problems in ProofNet (5.69) and the more self-contained problems in MiniF2F-test (5.21).
A.3 Prompt Templates
A.3.1 Decomposition Prompt
This appendix contains the complete prompt used to decompose informal mathematical statements into retrieval queries (the Decompose module). It is composed of two parts: a system prompt that defines the model’s expert persona and overall task, and a user prompt template that structures the specific input and desired output format.
A.3.2 Formalization Prompt
This appendix contains the complete prompt used to formalize informal mathematical statements into formal statements (the Formalize Theorems module). It is composed of two parts: a system prompt that defines the model’s expert persona and overall task, and a user prompt template that structures the specific input and desired output format.